# Ontological differentiation as a measure of semantic accuracy

### v2

@@ IV.6 Large-scale experiment design
    Random set pool: The pool of potential random concepts consisted of the 19,283 unique entries identified after the initial indexing and ground filtering stages (corresponding to the headwords in the ground-filtered dataset, described in Sec. IV.1).
@@ IV.7.1 Example: path comparison
[…]
@@
[…]
We then calculated the cumulative SOD score for each path using the formula $\text{SOD}_{C}(\mathcal{P})=\sum_{i=1}^{L}\text{SOD}(\text{origin},p_{i})$. The pairwise SOD scores between the origin (‘zebra’) and each subsequent node on the paths were computed (determining $\omega_{GOD}+1$ for each pair): SN path contributions:
@@
[…]
@@ V.1 Relationship between direct SOD and cosine similarity
The raw distributions of the two metrics differ substantially and vary across conditions. Detailed plots showing the distributions at each transformation stage (raw, log(1+x), normalized log(1+x)) for each condition are provided in the Supplementary Material (Figs. S1–S3). Consistently across all conditions, the raw SOD scores exhibit an extremely heavy right tail (see Figs. S1A, S2A, S3A), spanning many orders of magnitude and necessitating transformation for meaningful analysis. Raw cosine similarity values are bounded but also show skewness towards lower values (Figs. S1D, S2D, S3D). Given the extreme skewness, particularly of SOD, we applied a $\log(1+x)$ transformation to both metrics, a standard practice for handling such data. Figure 1 presents comparative box plots of the resulting log(1+SOD) and log(1+Cosine) distributions across the three conditions. These plots reveal differences in the central tendency and spread; for instance, the median log(1+Cosine) is notably lower for the targeted removal condition compared to the other two, while the log(1+SOD) distributions show variations primarily in their interquartile range and the extent of outliers. Kolmogorov-Smirnov tests confirmed that these log-transformed distributions differ significantly between all pairs of conditions (p $<$ 0.0001; see Sec. V.3), underscoring that the definition processing significantly impacts the resulting metric distributions.
@@
To quantify the monotonic relationship between the original metrics, we calculated the Spearman rank correlation coefficient ($\rho$) between the raw $\text{SOD}(s,t)$ scores and the raw $\cos(\mathbf{v}_{s},\mathbf{v}_{t})$ values for each condition. Table 2 summarizes these results. A clear trend emerges: the ground-filtered dataset exhibits a weak negative correlation ($\rho\approx-0.21$), which diminishes for the random removal dataset ($\rho\approx-0.10$) and becomes practically negligible for the targeted removal dataset ($\rho\approx-0.03$). While statistically significant due to the large sample sizes ($p<0.0001$ for all three conditions), even the strongest correlation observed ($-0.21$) is weak in practical terms. This indicates no meaningful rank-order relationship in the targeted removal case, and only a very slight tendency for higher raw SOD scores to correspond with lower raw cosine similarities in the less processed conditions.
@@
Ground-filtered $\approx 2.0\times 10^{6}$0.2069 ($p<0.0001$)
Random removal $\approx 2.0\times 10^{6}$0.0984 ($p<0.0001$)
Targeted removal $\approx 2.0\times 10^{6}$0.0305 ($p<0.0001$)
@@
Taken together, these analyses consistently demonstrate a weak relationship between direct SOD scores and RIM cosine similarity across all dataset conditions. This relationship, characterized generally by weak negative rank correlation (on raw data) and low percentile agreement (after log transformation), becomes progressively weaker, approaching statistical orthogonality, as definitions undergo more specific curation (from ground-filtered to targeted removal). This trend strongly suggests that SOD captures distinct aspects of semantic relatedness grounded in explicit definitional structure, which are largely independent of the network proximity information captured by RIM cosine similarity, particularly when applied to a more curated, semantically focused network like the targeted removal dataset.
@@
Table 4 presents the aggregated results comparing the $\text{SOD}_{C}$ scores for paths generated by SN versus those by SP between the same endpoints for each condition. A consistent and robust finding emerges across all three processing schemes: paths generated by SN exhibit lower $\text{SOD}_{C}$ scores than paths generated by SP in a significant majority of cases (59.82% for ground-filtered, 62.42% for random removal, and 58.22% for targeted removal). This primary observation holds regardless of the definition processing method applied. Statistical tests further confirm the significance of this dominance within each condition: $\chi^{2}$ goodness-of-fit tests strongly reject the hypothesis of equal likelihood for the three outcomes, namely: $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$, $\text{SOD}_{C}(\text{SN})=\text{SOD}_{C}(\text{SP})$, $\text{SOD}_{C}(\text{SN})>\text{SOD}_{C}(\text{SP})$ ($p<$ 0.0001 for all conditions). Furthermore, one-sided binomial tests confirm that the proportion of pairs where $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$ is significantly greater than 50% for each condition ($p<$ 0.0001). This provides strong evidence that SN consistently generates paths more definitionally coherent with the origin (as measured by lower $\text{SOD}_{C}$) compared to SP, demonstrating the robustness of using this cumulative SOD metric for validating semantic navigation across different corpus characteristics. This fundamental result is further tested and confirmed on more restrictive directed networks in the Supplemental Material, where we also control for the confounding effect of path length.
@@ V.3 Statistical comparison of metric distributions
Table 5: Kolmogorov-Smirnov (KS) 2-sample test results comparing log(1+SOD) and log(1+Cosine) distributions between conditions: ground-filtered (GF), random removal (RR), targeted removal (TR).
@@ VI Conclusions
Our investigation yielded two primary conclusions. First, we examined the relationship between direct pairwise SOD scores and RIM cosine similarity across the three network conditions. The results revealed that this relationship is sensitive to definition processing: a weak negative correlation in the ground-filtered network ($\rho\approx-0.21$) progressively diminished to become negligible in the targeted removal network ($\rho\approx-0.03$) (Table 2). This trend, supported by percentile agreement analyses (Table 3), establishes that SOD captures aspects of semantic relatedness grounded in explicit definitions that are largely orthogonal to embedding-based similarity, particularly as the definitional structure is refined. This confirms SOD offers a complementary perspective to standard similarity measures.
@@
In essence, Ontological Differentiation provides a formal method grounded in shared definitional components. Our comparative analysis across systematically varied definition sets validates SOD as a definition-based measure of semantic divergence, distinct from vector similarity. Most importantly, its consistent ability to identify SN paths as more definitionally coherent than SP paths across varied network structures demonstrates its robustness and utility as a tool for analyzing and validating semantic processes in lexical networks, offering insights independent of specific corpus construction details. This conclusion is not only supported by our main analysis on symmetric networks but is decisively reinforced by our extended analysis in the Supplemental Material. There, we demonstrate that SN’s superiority holds even on structurally constrained directed networks once the confounding effect of path length is properly controlled, confirming that the semantic coherence captured by SN is a fundamental advantage of the navigation strategy itself, independent of specific corpus construction or network topology.

### v3

@@ IV.6 Large-scale experiment design
    Complete pool: The pool of potential random concepts consisted of the 19,283 unique entries identified after the initial indexing and ground filtering stages (corresponding to the headwords in the ground-filtered dataset, described in Sec. IV.1).
@@ IV.7.1 Example: path comparison
$\mathcal{P}_{\rm SN}=(\text{`zebra', `animal', `move', `whirl', `circle'})$
@@
$\mathcal{Q}_{\rm SP}=(\text{`zebra', `looks', `full moon', `circle'})$
We then calculated the cumulative SOD score for each path using the formula $\text{SOD}_{C}(\mathcal{P})=\sum_{i=1}^{L}\text{SOD}(\text{origin},p_{i})$. The pairwise SOD scores between the origin (‘zebra’) and each subsequent node on the paths were computed (determining $\omega_{GOD}+1$ for each pair):
SN path contributions:
@@
In this example, the SN path exhibits a cumulative SOD score nearly an order of magnitude lower than the SP score. This stark difference is primarily driven by the extremely high pairwise $\text{SOD}(\text{zebra, full moon})=388098$ encountered along the SP route, signifying a major definitional divergence from the origin (‘zebra’) introduced by the ‘full moon’ node. While the SP path (‘zebra’ $\rightarrow$ ‘looks’ $\rightarrow$ ‘full moon’ $\rightarrow$ ‘circle’) is optimal based on structural edge weights (likely exploiting shared terms like ‘looks’), it appears intuitively less semantically direct than the SN path (‘zebra’ $\rightarrow$ ‘animal’ $\rightarrow$ ‘move’ $\rightarrow$ ‘whirl’ $\rightarrow$ ‘circle’).
@@ V.1 Relationship between direct SOD and cosine similarity
The raw distributions of the two metrics differ substantially and vary across conditions. Detailed plots showing the distributions at each transformation stage (raw, log(1+x), normalized log(1+x)) for each condition are provided in the Supplementary Material (Figs. S1–S3). Consistently across all conditions, the raw SOD scores exhibit an extremely heavy right tail (see Figs. S1A, S2A, S3A), spanning many orders of magnitude and necessitating transformation for meaningful analysis. Raw cosine similarity values are bounded but also show skewness towards lower values (Figs. S1D, S2D, S3D). Given the extreme skewness, particularly of SOD, we applied a $\log(1+x)$ transformation to both metrics, a standard practice for handling such data. Figure 1 presents comparative box plots of the resulting log(1+SOD) and log(1+Cosine) distributions across the three conditions. These plots reveal differences in the central tendency and spread; for instance, the median log(1+Cosine) is notably lower for the targeted removal condition compared to the other two, while the log(1+SOD) distributions show variations primarily in their interquartile range and the extent of outliers. Kolmogorov-Smirnov tests confirmed that these log-transformed distributions differ significantly between all pairs of conditions (p << 0.0001; see Sec. V.3), underscoring that the definition processing significantly impacts the resulting metric distributions.
@@
To quantify the monotonic relationship between the original metrics, we calculated the Spearman rank correlation coefficient ($\rho$) between the raw $\text{SOD}(s,t)$ scores and the raw $\cos(\mathbf{v}_{s},\mathbf{v}_{t})$ values for each condition. Table 2 summarizes these results. A clear trend emerges: the ground-filtered dataset exhibits a weak negative correlation ($\rho\approx-0.21$), which diminishes for the random removal dataset ($\rho\approx-0.10$) and becomes practically negligible for the targeted removal dataset ($\rho\approx-0.03$). While statistically significant due to the large sample sizes ($p<0.0001$ for all three conditions), even the strongest correlation observed (0.21-0.21) is weak in practical terms. This indicates no meaningful rank-order relationship in the targeted removal case, and only a very slight tendency for higher raw SOD scores to correspond with lower raw cosine similarities in the less processed conditions.
@@
Ground-filtered $\approx 2.0\times 10^{6}$ -0.2069 ($p<0.0001$)
Random removal $\approx 2.0\times 10^{6}$ -0.0984 ($p<0.0001$)
Targeted removal $\approx 2.0\times 10^{6}$ -0.0305 ($p<0.0001$)
@@
Conceptual comparison through illustrative examples.
To provide a more intuitive and qualitative understanding of this orthogonality, we performed a focused analysis on a curated set of nine concepts: ‘dog’, ‘moon’, ‘sea’, ‘doctor’, ‘war’, ‘hospital’, ‘police’, ‘soldier’, and ‘circus’. This set was specifically chosen because it contains pairs with a priori intuitive semantic relationships of varying degrees. This allows us to evaluate how well each metric’s similarity ranking aligns with human intuition. Before presenting the results, we can hypothesize a more nuanced hierarchy of relationships:
  • First-order relationships: Strong, direct connections such as (‘doctor’, ‘hospital’), (‘war’, ‘soldier’), and the physical/conceptual link (‘moon’, ‘sea’). A robust semantic metric should rank these highest.
  • Second-order relationships: Clear but less direct links, including functional associations like (‘soldier’, ‘police’), (‘dog’, ‘police’), (‘sea’, ‘police’), or thematic connections like (‘doctor’, ‘war’).
  • Third-order relationships: More abstract or indirect links, such as (‘war’, ‘hospital’) and (‘doctor’, ‘police’).
We computed the direct SOD score and cosine similarity for all 36 unique combinations of these words using the ground-filtered dataset. The top five and bottom five most similar pairs, as ranked by each metric, are presented in Table 4.
  Rank   SOD   Cosine
  1   (doctor, hospital)   (doctor, hospital)
  2   (war, soldier)   (moon, sea)
  3   (moon, sea)   (war, police)
  4   (sea, police)   (doctor, police)
  5   (doctor, war)   (hospital, police)
  32   (dog, soldier)   (war, circus)
  33   (dog, war)   (soldier, circus)
  34   (dog, circus)   (dog, circus)
  35   (soldier, circus)   (dog, soldier)
  36   (war, circus)   (dog, hospital)
Table 4: The five most similar (top) and five least similar (bottom) word pairs from a curated set of 36, as ranked by SOD and Cosine Similarity.
The rankings in Table 4 reveal the distinct nature of the metrics and their alignment with our intuitive hierarchy. The SOD rankings are remarkably consistent with our a priori expectations. Its top three pairs—(‘doctor’, ‘hospital’), (‘war’, ‘soldier’), and (‘moon’, ‘sea’)—are precisely the three we identified as having first-order relationships. Furthermore, its top five is rounded out by the strong second-order pairs (‘sea’, ‘police’) and (‘doctor’, ‘war’). This demonstrates SOD’s effectiveness in capturing structurally significant definitional links. In contrast, while cosine similarity also identifies (‘doctor’, ‘hospital’) and (‘moon’, ‘sea’) in its top two, it fails to recognize the salient ‘war-soldier’ connection, which does not appear in its top five. Instead, its rankings are heavily influenced by thematic proximity, with three of its top five pairs involving ‘police’ in combination with other civil or emergency-service concepts.
For pairs not captured in the top five, the results are more mixed and further highlight the trade-offs between the metrics. The third-order pair (‘war’, ‘hospital’) is ranked similarly by both but still better by SOD (14th by SOD, 16th by cosine). However, cosine similarity aligns better with some of our hypothesized second-order functional relationships, ranking (‘soldier’, ‘police’) 7th, whereas SOD ranks it a distant 25th. Similarly, (‘dog’, ‘police’) is ranked 24th by cosine versus 29th by SOD. This reinforces the idea that cosine is highly adept at capturing thematic and functional associations derived from contextual usage.
Looking at the bottom of the rankings, we observe a strong consensus. Both metrics agree that pairs involving ‘circus’ are semantically distant from the other concepts, and both rank the pair ‘dog-soldier’ as highly dissimilar. This concordance on clearly unrelated pairs validates that both metrics are behaving sensibly. The superior alignment of SOD’s top rankings with our intuitively structured hypothesis, however, highlights its strength as a measure of definition-grounded semantics. These illustrative examples substantiate our large-scale findings, confirming that SOD measures semantic proximity through definitional overlap, while cosine similarity measures it through contextual association.
@@
Taken together, our large-scale quantitative and focused qualitative analyses consistently demonstrate a weak relationship between direct SOD scores and RIM cosine similarity. The low rank correlation and percentile agreement across millions of pairs indicate statistical orthogonality, and our illustrative examples provide the conceptual grounding for this finding. The qualitative comparison revealed that SOD’s rankings align remarkably well with a priori definitional hierarchies, successfully identifying core relationships such as (‘war’, ‘soldier’) that are missed by cosine similarity. Conversely, cosine excels at identifying the thematic and contextual associations for pairs like (‘soldier’, ‘police’) and (‘dog’, ‘police’). This trend, which becomes more pronounced as definitions are curated, confirms that SOD captures a distinct aspect of semantic relatedness grounded in explicit definitional structure, which is largely independent of the network proximity and contextual information captured by RIM cosine similarity.
@@
Table 5 presents the aggregated results comparing the $\text{SOD}_{C}$ scores for paths generated by SN versus those by SP between the same endpoints for each condition. A consistent and robust finding emerges across all three processing schemes: paths generated by SN exhibit lower $\text{SOD}_{C}$ scores than paths generated by SP in a significant majority of cases (59.82% for ground-filtered, 62.42% for random removal, and 58.22% for targeted removal). This primary observation holds regardless of the definition processing method applied. Statistical tests further confirm the significance of this dominance within each condition: $\chi^{2}$ goodness-of-fit tests strongly reject the hypothesis of equal likelihood for the three outcomes, namely: $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$, $\text{SOD}_{C}(\text{SN})=\text{SOD}_{C}(\text{SP})$, $\text{SOD}_{C}(\text{SN})>\text{SOD}_{C}(\text{SP})$ ($p<$ 0.0001 for all conditions). Furthermore, one-sided binomial tests confirm that the proportion of pairs where $\text{SOD}_{C}(\text{SN})<\text{SOD}_{C}(\text{SP})$ is significantly greater than 50% for each condition ($p<$ 0.0001). This provides strong evidence that SN consistently generates paths more definitionally coherent with the origin (as measured by lower $\text{SOD}_{C}$) compared to SP, demonstrating the robustness of using this cumulative SOD metric for validating semantic navigation across different corpus characteristics. This fundamental result is further tested and confirmed on more restrictive directed networks in the SM [19], where we also control for the confounding effect of path length.
@@ V.3 Statistical comparison of metric distributions
Table 6: Kolmogorov-Smirnov (KS) 2-sample test results comparing log(1+SOD) and log(1+Cosine) distributions between conditions: ground-filtered (GF), random removal ($R$), targeted removal (TR).
@@ VI Conclusions
Our investigation yielded several key conclusions. First, we examined the relationship between direct pairwise SOD scores and RIM cosine similarity. The results revealed that this relationship is sensitive to definition processing: a weak negative correlation in the ground-filtered network ($\rho\approx-0.21$) progressively diminished to become negligible in the targeted removal network ($\rho\approx-0.03$) (Table 2). This trend, supported by percentile agreement analyses (Table 3), establishes that SOD is largely orthogonal to embedding-based similarity. This statistical independence is grounded in a fundamental conceptual distinction, as highlighted by our illustrative examples (Table 4). The analysis revealed that SOD excels at capturing explicit definitional hierarchies, successfully identifying core relationships such as (‘war’, ‘soldier’) that are missed by cosine similarity’s top rankings. Conversely, cosine similarity proved more adept at identifying thematic and contextual associations, such as (‘soldier’, ‘police’). This confirms that SOD offers a complementary, definition-grounded perspective on semantic relatedness that is distinct from existing vector-space and gloss-overlap methods.
@@
This finding—that SN paths consistently exhibit lower cumulative SOD scores than SP paths—offers a key insight into the nature of navigation in semantic networks. This result highlights a fundamental distinction between two navigation philosophies: structural economy versus semantic coherence. The Shortest Path (SP) algorithm, by design, optimizes for structural economy, identifying the most direct path available through the explicit definitional links. In contrast, Semantic Navigation (SN) uses vector embeddings that encode high-order, network-wide relationships to maintain coherence with the target. The fact that SOD, a metric grounded purely in the dictionary’s explicit definitional structure, consistently favors SN paths is a non-trivial result. It provides strong quantitative evidence that the statistical patterns learned by the RIM embedding method successfully capture the underlying semantic coherence of the lexicon. In other words, SN’s greedy, vector-guided strategy produces trajectories that are not only statistically plausible but are also more faithful to the formal meaning encoded in dictionary definitions. This validates the use of such embeddings for generating semantically meaningful pathways and demonstrates that SOD can serve as an effective, independent benchmark for measuring this alignment between statistical and symbolic representations of knowledge.
Beyond its role as an evaluative metric, SOD can be regarded as a constructive heuristic for semantic navigation itself. In the Section II of the SM [19], we introduce a proof-of-concept for semantic navigation based on SOD (which we term Ontological Navigation, ON), based on greedily selecting paths by minimizing the direct SOD score to the target. Qualitative analysis of the resulting paths (see Supplemental Table S1) shows they are often qualitatively different from those of SN and follow a more intuitive, definition-based logic, exemplified by the path ‘dog’ $\rightarrow$ ‘feist’ $\rightarrow$ ‘belligerent’ $\rightarrow$ ‘war’. This demonstrates that SOD is not merely a passive validation tool but can serve as an active and principled heuristic, opening a new avenue for purely symbolic semantic navigation.
In essence, Ontological Differentiation provides a formal method grounded in shared definitional components. Our comparative analysis across systematically varied definition sets validates SOD as a definition-based measure of semantic divergence, distinct from vector similarity. Most importantly, its consistent ability to identify SN paths as more definitionally coherent than SP paths across varied network structures demonstrates its robustness and utility as a tool for analyzing and validating semantic processes in lexical networks, offering insights independent of specific corpus construction details. Furthermore, this work demonstrates its dual potential: beyond its power as an analytical tool, its promise as a constructive heuristic is introduced through our proof-of-concept Ontological Navigation. Our analysis showing that SOD is coherently sensitive to polysemy suggests promising future extensions for this navigation, potentially allowing for context-aware pathfinding. The conclusions drawn from our main analysis on symmetric networks are decisively reinforced by our extended analysis in the SM [19]. There, we show that the superior coherence of SN paths holds even on structurally constrained directed networks once the confounding effect of path length is properly controlled, thereby validating SOD’s capacity to serve as a robust evaluation tool whose findings are independent of specific corpus construction or network topology. The principled nature of the framework, supported by this demonstrated robustness, also points to its generalizability beyond the linguistic domain to mathematical and physical ontologies, a subject of ongoing investigation.